OpenStudy (anonymous):

What is this equation in standard form? y^2+4y-8x+4=0 I believe the standard form for a hyperbola is (x-h)^2/a^2 - (y-k)^2/b^2=1 I just don't know what terms to put where.

OpenStudy (anonymous):

@ikram002p Do you know how to do this?

OpenStudy (ikram002p):

y^2+4y-8x+4=0 y^2+4y+4-8x=0 (y+2)^2-8x=0 x=(y+2)^2/8 so its parabola

OpenStudy (anonymous):

Argh. That's probably why I couldn't make it work in the first place. Thanks!

OpenStudy (ikram002p):

wlc